OpenStudy (anonymous):

Evaluate the logorithm log(6)1/36

OpenStudy (owlcoffee):

I would recommend effectuating the corresponding exponential transformations and trying to form a base which respects the one in the logarithm. Since the level is labeled "high" I suppose you have no problem with that, so, just stating a possible method to evaluate the logarithm without calculator: \[36=6^2\] \[\log_{6} \frac{ 1 }{ 36 } \iff \log_{6} \frac{ 1 }{ 6^2 }\] With that transformed, I suggest you use the exponential property of negative exponents and bases on denominator, to posterior use the logarithmic property of exponential logarithmand. Here they are: \[\frac{ 1 }{ a^m }=a ^{-m}\] \[\log_{a} b^n \iff (n)\log_{a} b\]

OpenStudy (anonymous):

So 6 is a, m is 2, b is 1/36? And then I just plug them in? @Owlcoffee

OpenStudy (owlcoffee):

remember I transformed the 36 into 6^2.

OpenStudy (anonymous):

ok so it would be 1/6^2 = 6^-2? @Owlcoffee

OpenStudy (owlcoffee):

That is correct.

OpenStudy (anonymous):

How do I solve that or is that the answer? @Owlcoffee

OpenStudy (owlcoffee):

So far you have \[\log_{6} 6^{-2}\] You can use the other property of logarithms I told you: \[\log_{a} b^n \iff (n)\log_{a} b\]

OpenStudy (anonymous):

Which number is n? @Owlcoffee

OpenStudy (owlcoffee):

Observe the structure "n" is the exponent above "6"

OpenStudy (anonymous):

ok so -2. So it would be log(6)36^-2 <----->(-2)log(6) 36 @Owlcoffee

OpenStudy (owlcoffee):

That is not correct, don't forget that it has to be strictly the 6^-2. 6... not 36.

OpenStudy (anonymous):

oh jeez right. Ok log(6)6^-2 <---->(-2)log(6)6 @Owlcoffee

OpenStudy (owlcoffee):

That is correct. Now, what is \(\log_{6} 6\) equal to?

OpenStudy (anonymous):

-2? @Owlcoffee

OpenStudy (owlcoffee):

That's incorrect. To what exponent do you have to raise "6" in order to obtain as a result "6"?

OpenStudy (anonymous):

1? I'm sorry i'm just really bad at math. it might take me a while....

OpenStudy (owlcoffee):

That is correct \[\log_{6} 6=1\]

OpenStudy (anonymous):

Yes! THANK YOU!

OpenStudy (owlcoffee):

So, what would be the result of this? \[(-2)\log_{6} 6\]

OpenStudy (anonymous):

-2 right?

OpenStudy (owlcoffee):

Correct.

OpenStudy (anonymous):

Yay! I really appreciate your help amd patience. Like this helped sooo much.

OpenStudy (owlcoffee):

No problem, that's why we're helpers here.